OpenStudy (anonymous):

if f(x)=x^2+3x+5, what is f(a+h)?

OpenStudy (anonymous):

put x=a+h

OpenStudy (nurali):

f(x)=x^2+3x+5 f(a+h)=(a+h)^2+3(a+h)+5 f(a+h)=a^2+2ah+h^2+3a+3h+5 f(a+h)=a^2+2ah+3a+3h+h^2+5